# Coulomb branch surgery: Holonomy saddles, S-folds and discrete symmetry gaugings

### v1

@@ 1.1 Automorphisms of the SW geometry
In this paper, we extend the applications of SW geometry through a careful analysis of its symmetries. We direct our attention to 4d rank-one $\mathcal{N}=2$ theories, which also include Kaluza–Klein (KK) theories. There, the total space of the SW curve fibred over the complex one-dimensional Coulomb branch, or $U$-plane, can be identified as a rational elliptic surface (RES) Caorsi:2018ahl ; Closset:2021lhd . In many instances, quotients of a RES by subgroups of the full automorphism group ${\rm Aut}(\mathcal{S})$ lead to other well-defined RES. As a result, a link between seemingly distinct SW geometries is realised. Rather remarkably, these surgeries lead to deep physical insights, going beyond discrete symmetry gaugings, contrary to naive expectations.
@@ 1.3 Outline
Several details and definitions of elliptic surfaces as well as modular forms are given in Appendix A. In Appendix B, we provide a more general description of how $U$-plane foldings can be obtained from the perspective of fundamental domains. Appendix C discusses further aspects of non-cyclic Coulomb branch symmetries by analysing several examples. Appendix D finally lists some explicit expressions for the Seiberg–Witten curves. We also include a Mathematica notebook with explicit curves for all $D_{S^{1}}E_{n}$ theories encurves .
@@ 2.1 Introducing the $U$-plane
Finally, let us comment on the higher-form symmetries of the theories described above. The candidate theories that have 1-form symmetries are those which allow a gauge theory description where the matter fields do not transform under the centre of the gauge group Gaiotto:2014kfa . The first such example is the pure 4d $\text{SU}(2)$ theory, which has a $\Gamma^{(1)}=\mathbb{Z}_{2}$ 1-form symmetry. For future reference, let us mention that this theory also has a mixed anomaly between the 1-form symmetry and the $\mathbb{Z}_{8}\subset{\rm U}(1)_{\mathtt{r}}$ 0-form symmetry, originating from the classical $R$-symmetry and broken by an ABJ anomaly Cordova:2018acb . Another example of a purely 4d theory having a 1-form symmetry is the $\mathcal{N}=2^{*}$ theory, which is the $\mathcal{N}=2$ preserving mass deformation of $\mathcal{N}=4$ SYM.
@@ 4.2 The massless cases
Fortunately, it was shown in DelMonte:2021ytz that for this particular quiver there exists a chamber where the BPS spectrum decomposes into two copies of the 4d ${\rm SU}(2)$ massless $N_{f}=1$ theory, together with a tower of KK states. The corresponding 4d sub-quivers are coloured in the above diagram.
@@
The relevant $\mathbb{Z}_{2}$ symmetric $D_{S^{1}}E_{7}$ quiver is given by: [^16] This is mutation equivalent to other known $dP_{7}$ quivers Wijnholt:2002qz .
@@
As before, we use the block notation. We again expect that there is a chamber where the BPS spectrum of the KK theory decomposes into two copies of the massless $N_{f}=3$ sub-quivers. To the best of our knowledge, this has not been discussed in the literature.
@@ 4.3.1 The 5d–4d dictionary
where $u_{\mathcal{N}=2^{*}}$ is the 4d $\mathcal{N}=2^{*}$ Coulomb branch parameter, $U_{\text{M-string}}$ parametrises the M-string curve, and $\beta$ is the radius of the $S^{1}$. Meanwhile, Weierstrass-$\wp$ is the Weierstraß function. A BPS quiver for the 6d M-string was proposed in Closset:2023pmc , by making use of this map.
@@ 5.2 Examples
The ‘magnetic’ theory would be obtained by gauging both the 0 and 1-form symmetries of the KK theory, which is equivalent to gauging the whole 1-form symmetry for the 5d SCFT. However, our proposal will not apply in this case.
@@ Discussion.
Let us also comment on a possible connection to 2-groups. The semi-direct product ${\rm Aut}(\mathcal{S})$ is induced by a homomorphism $\varphi:{\rm Aut}_{\sigma}(\mathcal{S})\to{\rm Aut}({\rm MW}(\mathcal{S}))$, as defined precisely in (A.11) in Appendix A.2. Meanwhile, 0-form and 1-form global symmetries can combine into 2-group global symmetries Apruzzi:2021vcu ; Cordova:2018cvg ; Benini:2018reh ; Albertini:2020mdx . Since 0-form symmetries can act on the 1-form charges, the symmetry defect demands an action $\rho:\Gamma^{(0)}\to\text{Aut}(\Gamma^{(1)})$. Together with the Postnikov class $[\beta]\in H_{\rho}^{3}(B\Gamma^{(0)},\Gamma^{(1)})$, this action defines a general 2-group $\mathbb{G}=(\Gamma^{(0)},\Gamma^{(1)},\rho,[\beta])$ Benini:2018reh . Our discussion elucidates the way in which the homomorphism $\varphi$ descends to the action $\rho$. It would be interesting to connect these ideas to theories that indeed admit 2-groups, as well as non-invertible symmetries Kaidi:2021xfk ; Apruzzi:2021vcu ; Cordova:2018cvg ; Benini:2018reh ; Albertini:2020mdx ; Closset:2021lhd . See also Arias-Tamargo:2023duo for a related approach.
@@ Acknowledgements.
We are very grateful to Johannes Aspman, Cyril Closset and Jan Manschot for many interesting discussions, feedback and comments on the draft. EF is supported by the EPSRC grant “Local Mirror Symmetry and Five-dimensional Field Theory”. HM was supported by a Royal Society Research Grant for Research Fellows for the most part of this work.
@@ A.2 Automorphisms of elliptic surfaces
Consider a rational elliptic surface $\mathcal{S}$. The rational sections $\beta:\mathcal{S}\to\mathbb{P}^{1}$ of $\mathcal{S}$ form the Mordell–Weil group ${\rm MW}(\mathcal{S})$, which is a subgroup of the group of automorphisms of $\mathcal{S}$, ${\rm Aut}(\mathcal{S})$ KARAYAYLA20121 ; Karayayla+2014+1772+1795 ; Bouchard:2007mf . It can be shown (see KARAYAYLA20121 ) that the automorphism group of $\mathcal{S}$ is isomorphic to the semi-direct product [^22] Recall that for two groups $G$ and $H$, a group homomorphism $\varphi:G\to\rm Aut(H)$ defines a semi-direct product $H\rtimes_{\varphi}G\subset H\times G$ with the multiplication $(h_{1},g_{1})(h_{2},g_{2})\coloneqq(h_{1}\varphi(g_{1})(h_{2}),g_{1}g_{2})$. For $(h,g)\in H\rtimes_{\varphi}G$, the inverse is found as $(\varphi(g^{-1})(h^{-1}),g^{-1})$.
@@
The automorphism group preserving the zero section is related to the automorphism group on the base $\mathbb{P}^{1}$ in a simple way: For every automorphism $\tau$ of $\mathcal{S}$, there is an induced automorphism $\tau_{\mathcal{S}}$ on $\mathbb{P}^{1}$ such that the obvious corresponding diagram with rational sections $\beta:\mathcal{S}\to\mathbb{P}^{1}$ is commutative. There is another group homomorphism
@@ A.3 Modular forms and congruence subgroups
and are related by conjugation. The principal congruence subgroup $\Gamma(n)$ is the subgroup of $\text{SL}(2,\mathbb{Z})\ni A$ defined by elements $A\equiv\mathbbm{1}\mod n$. A subgroup $\Gamma$ of $\text{SL}(2,\mathbb{Z})$ is called a congruence subgroup if it contains $\Gamma(n)$ for some $n\in\mathbb{N}$.
@@
Let us also study the projective indices of several subgroups of $\text{PSL}(2,\mathbb{Z})$, [^24] See for instance Diamond ; koblitz1993
@@
$\text{SL}(2,\mathbb{Z})/\Gamma(n)\cong\text{SL}(2,\mathbb{Z}_{n})\leavevmode\nobreak\ .$ (A.28)
@@
The order of $\Gamma/\Gamma(n)$ matches of course with the index $[\Gamma(1):\Gamma(n)]$. It is clear that $\text{SL}(2,\mathbb{Z})/\Gamma(n)$ for larger $n$ cannot be isomorphic to $S_{m}$ or $A_{m}$, since $|S_{m}|\sim m!$, while $[\Gamma(1):\Gamma(n)]\sim n^{3}$.
@@ Appendix B Foldings of fundamental domains
An approach to define an analogue group for pairs of non-normal subgroups is the following. Let $\Gamma(1)=\text{PSL}(2,\mathbb{Z})$, and let $\Gamma$ be any finite index subgroup of $\Gamma(1)$. Then we can decompose $\Gamma(1)$ into a finite union
@@ Appendix C Non-cyclic symmetries
The Coulomb branch automorphisms induced by automorphisms of the SW geometry are discrete subgroups of $\text{Aut}(\mathbb{P}^{1})\cong\text{PGL}(2,\mathbb{Z})$ of order less than or equal to 12.313131Automorphisms on $\mathbb{P}^{1}$ are bijective conformal maps. They are: the cyclic groups $\mathbb{Z}_{n}$ with $n\leq 12$, the dihedral groups [^32] For future reference, recall that the dihedral group $D_{n}$ with $2n$ elements is the symmetry group of the regular $n$-gon and has a presentation $D_{n}=\langle r,s\,|\,r^{n}=s^{2}=(sr)^{2}=e\rangle\leavevmode\nobreak\ ,$ (C.1) such that it is isomorphic to $D_{n}\cong\mathbb{Z}_{n}\rtimes\mathbb{Z}_{2}$. $D_{n}$ (of order $2n$) with $n=2,3,4,6$, and the tetrahedral group $A_{4}$. The family of RES supporting the non-cyclic $\text{Aut}_{\mathcal{S}}(\mathbb{P}^{1})$ is classified by their singular fibres, as given in (KARAYAYLA20121, , Proposition 4.2.3).
@@ Appendix D Seiberg–Witten curves
See also footnote 18 for an important comment on the normalisation for $N_{f}=3$.

### v2

@@ 1.1 Automorphisms of the SW geometry
In this paper, we extend the applications of SW geometry through a careful analysis of its symmetries. We direct our attention to 4d rank-one $\mathcal{N}=2$ theories, which also include Kaluza–Klein (KK) theories. There, the total space of the SW curve fibred over the complex one-dimensional Coulomb branch, or $U$-plane, can be identified as a rational elliptic surface (RES) Caorsi:2018ahl ; Closset:2021lhd . In many instances, quotients of a RES by subgroups of the full automorphism group ${\rm Aut}(\mathcal{S})$ lead to other well-defined RES. As a result, a link between seemingly distinct SW geometries is realised. Rather remarkably, these surgeries [^1] This term should not be mistaken for the one used in surgery theory in the field of algebraic topology. lead to deep physical insights, going beyond discrete symmetry gaugings, contrary to naive expectations.
@@ 1.3 Outline
Several details and definitions of elliptic surfaces as well as modular forms are given in Appendix A. In Appendix B, we provide a more general description of how $U$-plane foldings can be obtained from the perspective of fundamental domains. Appendix C discusses further aspects of non-cyclic Coulomb branch symmetries by analysing several examples. Appendix D finally lists some explicit expressions for the Seiberg–Witten curves. We also include a list of Mathematica notebooks with explicit curves for all $D_{S^{1}}E_{n}$ theories, as well as various other routines used throughout this work in the GitHub repository encurves .
@@ 2.1 Introducing the $U$-plane
Finally, let us comment on the higher-form symmetries of the theories described above (see also DelZotto:2015isa ; Gukov:2020btk ; Bhardwaj:2021pfz ; Bhardwaj:2021mzl and references therein). The candidate theories that have 1-form symmetries are those which allow a gauge theory description where the matter fields do not transform under the centre of the gauge group Gaiotto:2014kfa . The first such example is the pure 4d $\text{SU}(2)$ theory, which has a $\Gamma^{(1)}=\mathbb{Z}_{2}$ 1-form symmetry. For future reference, let us mention that this theory also has a mixed anomaly between the 1-form symmetry and the $\mathbb{Z}_{8}\subset{\rm U}(1)_{\mathtt{r}}$ 0-form symmetry, originating from the classical $R$-symmetry and broken by an ABJ anomaly Cordova:2018acb . Another example of a purely 4d theory having a 1-form symmetry is the $\mathcal{N}=2^{*}$ theory, which is the $\mathcal{N}=2$ preserving mass deformation of $\mathcal{N}=4$ SYM.
@@ 3 Folding the $U$-plane
The recent work of Bershtein2021 on foldings of Painlevé equations leads to similar results to our CB folding analysis. These similarities are due to the close link between 5d BPS quivers and $q$-deformed Painlevé equations, recently studied in e.g.  Bonelli:2020dcp ; Bonelli:2016qwg ; Bonelli:2017gdk .
@@ 4.2 The massless cases
Fortunately, it was shown in DelMonte:2021ytz ; Bonelli:2020dcp that for this particular quiver there exists a chamber where the BPS spectrum decomposes into two copies of the 4d ${\rm SU}(2)$ massless $N_{f}=1$ theory, together with a tower of KK states. The corresponding 4d sub-quivers are coloured in the above diagram.
@@
The relevant $\mathbb{Z}_{2}$ symmetric $D_{S^{1}}E_{7}$ quiver and a basis of BPS state can be found by starting with a modular configuration of the theory. The best example is the configuration $(I_{2};I_{8},2I_{1})$, whose monodromy group is conjugate to $\Gamma^{0}(8)$. A possible fundamental domain for this monodromy group will then have a width 2 cusp at $\tau=i\infty$, as well as:
$I_{1}\text{ at }\tau=-{1\over 2}\leavevmode\nobreak\ ,\qquad I_{8}\text{ at }\tau=0\leavevmode\nobreak\ ,\qquad I_{1}\text{ at }\tau={1\over 2}\leavevmode\nobreak\ .$ (4.25)
Following the prescription outlined in Magureanu:2022qym , these generate the light BPS states $(-2,-1)$, $8(1,0)$ and $(-2,1)$, respectively. Mutating on 4 of the monopole states leads to the new basis:
$\gamma_{1}\,:$ $\quad(2,-1)\leavevmode\nobreak\ ,$ $\qquad\qquad\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5}\,:$ $\quad(1,0)\leavevmode\nobreak\ ,$ (4.26)
$\gamma_{6}$ $\quad(-2,1)\leavevmode\nobreak\ ,$ $\qquad\qquad\gamma_{7},\gamma_{8},\gamma_{9},\gamma_{10}\,:$ $\quad(-1,0)\leavevmode\nobreak\ ,$
with the corresponding BPS quiver given by: [^17] This is mutation equivalent to other known $dP_{7}$ quivers Wijnholt:2002qz .
@@
As before, we use the block notation. We again expect that there is a chamber where the full BPS spectrum of the KK theory decomposes into two copies of the massless $N_{f}=3$ subquivers. To the best of our knowledge, this has not been discussed in the literature. As an additional argument towards this claim, we notice that the basis of light BPS states (4.26) satisfies this decomposition – see e.g. Bilal:1996sk for details on the BPS spectrum of the massless 4d SU(2) $N_{f}=3$ quiver.
@@ 4.3.1 The 5d–4d dictionary
where $u_{\mathcal{N}=2^{*}}$ is the 4d $\mathcal{N}=2^{*}$ Coulomb branch parameter, $U_{\text{M-string}}$ parametrises the M-string curve, and $\beta$ is the radius of the $S^{1}$. Meanwhile, Weierstrass-$\wp$ is the Weierstraß function. A BPS quiver for the 6d M-string was proposed in Closset:2023pmc , by making use of this map, which we also show here for completeness:
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{6}$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{5}$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{1}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{4}$$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{2}$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{3}$ (4.38)
In the above diagram, we also indicate the 4d $\mathcal{N}=2^{\ast}$ subquivers. Note that the proposed M-string quiver also inherits the $\mathbb{Z}_{3}$ symmetry of the 4d $\mathcal{N}=2^{\ast}$ subquivers Argyres:2016yzz , which is present for particular values of the UV coupling. [^21] We thank the anonymous reviewer for pointing this out. This extends to a $\mathbb{Z}_{6}$ symmetry for the 6d quiver, which is identical to the $\mathbb{Z}_{6}$ symmetry of the $E_{3}$ quiver (4.16). Importantly, this $\mathbb{Z}_{6}$ symmetry only exists for the ‘relative’ form of the M-string theory, i.e. for the SW curve with singular bulk fibres $(6I_{2})$.
@@ 5.2 Examples
The ‘magnetic’ theory would be obtained by gauging both the 0 and 1-form symmetries of the KK theory, which is equivalent to gauging the whole 1-form symmetry for the 5d SCFT. However, our proposal does not apply in this case, since there is no 0-form symmetry in the magnetic version; instead, the gauging of the 0-form symmetry leads to a 2-form symmetry, and the (absence of a) mixed anomaly between the resulting 2-form symmetry and the magnetic 1-form symmetry does not appear to be visible at the level of the SW geometry.
@@ Discussion.
Let us also comment on a possible connection to 2-groups. The semi-direct product ${\rm Aut}(\mathcal{S})$ is induced by a homomorphism $\varphi:{\rm Aut}_{\sigma}(\mathcal{S})\to{\rm Aut}({\rm MW}(\mathcal{S}))$, as defined precisely in (A.11) in Appendix A.2. Meanwhile, 0-form and 1-form global symmetries can combine into 2-group global symmetries Apruzzi:2021vcu ; Cordova:2018cvg ; Benini:2018reh ; Albertini:2020mdx ; DelZotto:2022joo . Since 0-form symmetries can act on the 1-form charges, the symmetry defect demands an action $\rho:\Gamma^{(0)}\to\text{Aut}(\Gamma^{(1)})$. Together with the Postnikov class $[\beta]\in H_{\rho}^{3}(B\Gamma^{(0)},\Gamma^{(1)})$, this action defines a general 2-group $\mathbb{G}=(\Gamma^{(0)},\Gamma^{(1)},\rho,[\beta])$ Benini:2018reh . Our discussion elucidates the way in which the homomorphism $\varphi$ descends to the action $\rho$. It would be interesting to connect these ideas to theories that indeed admit 2-groups, as well as non-invertible symmetries Kaidi:2021xfk ; Apruzzi:2021vcu ; Cordova:2018cvg ; Benini:2018reh ; Albertini:2020mdx ; Closset:2021lhd . See also Arias-Tamargo:2023duo for a related approach.
@@ Acknowledgements.
We are very grateful to Johannes Aspman, Cyril Closset and Jan Manschot for many interesting discussions, feedback and comments on the draft. We further thank Fabrizio Del Monte for discussions. EF is supported by the EPSRC grant “Local Mirror Symmetry and Five-dimensional Field Theory”. HM was supported by a Royal Society Research Grant for Research Fellows for the most part of this work.
@@ A.2 Automorphisms of elliptic surfaces
Consider a rational elliptic surface $\mathcal{S}$. The rational sections of $\beta:\mathcal{S}\to\mathbb{P}^{1}$ form the Mordell–Weil group ${\rm MW}(\mathcal{S})$, which is a subgroup of the group of automorphisms of $\mathcal{S}$, ${\rm Aut}(\mathcal{S})$ KARAYAYLA20121 ; Karayayla+2014+1772+1795 ; Bouchard:2007mf . It can be shown (see KARAYAYLA20121 ) that the automorphism group of $\mathcal{S}$ is isomorphic to the semi-direct product [^24] Recall that for two groups $G$ and $H$, a group homomorphism $\varphi:G\to\rm Aut(H)$ defines a semi-direct product $H\rtimes_{\varphi}G\subset H\times G$ with the multiplication $(h_{1},g_{1})(h_{2},g_{2})\coloneqq(h_{1}\varphi(g_{1})(h_{2}),g_{1}g_{2})$. For $(h,g)\in H\rtimes_{\varphi}G$, the inverse is found as $(\varphi(g^{-1})(h^{-1}),g^{-1})$.
@@
The automorphism group preserving the zero section is related to the automorphism group on the base $\mathbb{P}^{1}$ in a simple way: For every automorphism $\tau$ of $\mathcal{S}$, there is an induced automorphism $\tau_{\mathcal{S}}$ on $\mathbb{P}^{1}$ such that the obvious corresponding diagram with rational sections of $\mathcal{S}$ is commutative. There is another group homomorphism
@@ A.3 Modular forms and congruence subgroups
and are related by conjugation. The principal congruence subgroup $\Gamma(n)$ is the subgroup of $\text{SL}(2,\mathbb{Z})\ni A$ defined by elements $A\equiv\mathbbm{1}\mod n$. Note, in particular, that $\Gamma(1)=\text{SL}(2,\mathbb{Z})$. A subgroup $\Gamma$ of $\text{SL}(2,\mathbb{Z})$ is called a congruence subgroup if it contains $\Gamma(n)$ for some $n\in\mathbb{N}$.
@@
Let us also study the projective indices of several subgroups of $\text{PSL}(2,\mathbb{Z})$, [^26] See for instance Diamond ; koblitz1993 .
@@
$\Gamma(1)/\Gamma(n)\cong\text{SL}(2,\mathbb{Z}_{n})\leavevmode\nobreak\ .$ (A.28)
@@
The order of $\Gamma(1)/\Gamma(n)$ matches of course with the index $[\Gamma(1):\Gamma(n)]$. It is clear that $\Gamma(1)/\Gamma(n)$ for larger $n$ cannot be isomorphic to $S_{m}$ or $A_{m}$, since $|S_{m}|\sim m!$, while $[\Gamma(1):\Gamma(n)]\sim n^{3}$.
@@ Appendix B Foldings of fundamental domains
An approach to define an analogue group for pairs of non-normal subgroups is the following. Let $\Gamma$ be any finite index subgroup of $\Gamma(1)$. Then we can decompose $\Gamma(1)$ into a finite union
@@ Appendix C Non-cyclic symmetries
The Coulomb branch automorphisms induced by automorphisms of the SW geometry are discrete subgroups of $\text{Aut}(\mathbb{P}^{1})\cong\text{PGL}(2,\mathbb{Z})$ of order less than or equal to 12.333333Automorphisms on $\mathbb{P}^{1}$ are bijective conformal maps. They are: the cyclic groups $\mathbb{Z}_{n}$ with $n\leq 12$, the dihedral groups [^34] For future reference, recall that the dihedral group $D_{n}$ with $2n$ elements is the symmetry group of the regular $n$-gon and has a presentation $D_{n}=\langle r,s\,|\,r^{n}=s^{2}=(sr)^{2}=e\rangle\leavevmode\nobreak\ ,$ (C.1) such that it is isomorphic to $D_{n}\cong\mathbb{Z}_{n}\rtimes\mathbb{Z}_{2}$. $D_{n}$ (of order $2n$) with $n=2,3,4,6$, and the tetrahedral group $A_{4}$. The family of RES supporting the non-cyclic $\text{Aut}_{\mathcal{S}}(\mathbb{P}^{1})$ is classified by their singular fibres, as given in (KARAYAYLA20121, , Proposition 4.2.3).
@@ Appendix D Seiberg–Witten curves
See also footnote 19 for an important comment on the normalisation for $N_{f}=3$.